Q:-Let a point P(h,k) satisfies the equation x3+y3+3xy=1 , then the locus of point P is a


1.Straight line


2.Circle


3.Parabola


4.Hyperbola


please xplain in details

Grade:12

1 Answers

Badiuddin askIITians.ismu Expert
14 years ago

Dear Tapasranjan Das

x3+y3+3xy-1=0

(x+y)3 -3xy(x+y) +3xy -1 =0

(x+y)3 -13 -3xy(x+y-1) =0

(x+y-1)[(x+y)2 +1 +(x+y)] --3xy(x+y-1) =0

(x+y-1) [(x+y)2 +1 +(x+y) -3xy]=0

(x+y-1) [x2+y2 +1 +x+y -xy]=0

 x+y -1 =0 rpresent a straight line

 and for x2+y2 +1 +x+y -xy]=0  check value of  Δ

Δ will also come zero so it represent a pair of straight line

Please feel free to post as many doubts on our discussion forum as you can.
If you find any question Difficult to understand - post it here and we will get you the answer and detailed
solution very  quickly.
 We are all IITians and here to help you in your IIT JEE  & AIEEE preparation.

 All the best.
 
Regards,
Askiitians Experts
Badiuddin